OpenStudy (anonymous):

using f(t)=Yob^t, 1.8 billions sales in 2004 and 8 billion in 2008. t=o correspond to to the year 2004. Find an expontential model for this data?

OpenStudy (amistre64):

8 = 1.8 b^4 8/1.8 = b^4 4rt(8/1.8) = b

OpenStudy (amistre64):

it looks good, but is it...

OpenStudy (anonymous):

one sec

OpenStudy (amistre64):

b = 1.452 then

OpenStudy (anonymous):

yes thanks. i had switched the x any for one point. thanks